Suppose you are trying to construct a periodic table in an alternate universe. You discover element "X", which has two isotopes. Isotope #1 has a mass of 14.02 amu and an abundance of 29.5%. Isotope #2 has a mass of 15.02 amu. What atomic mass would you put on the periodic table for element "X"?

Please show all working out and neat hand writing please. Thank you.
Answer #1

Isotope 1:

Mass = 14.02 amu

Abundance = 29.5%

Isotope 2:

Mass = 15.02 amu

Abundance = 100 – 29.5 = 70.5%

Atomic mass of element X = (14.02 x 29.5)/100 + (15.02 x 70.5)/100 = (4.1359 + 10.5891) amu = 14.725 amu